OpenStudy (anonymous):

having a hard time with sin, cos, and tan with that zero like thing next to it (I don't remember the name). Help?

OpenStudy (anonymous):

\(\sin(\theta)\) ?

OpenStudy (anonymous):

Yes, that.

OpenStudy (anonymous):

I'm needing to find the tan, sin, and cos for that triangle and it's online course so i don't really have any help.

OpenStudy (anonymous):

sine is opposite over hypotenuse you find the opposite side by pythagoras

OpenStudy (anonymous):

lets call that side \(b\) then pythagoras tells you that \[10^2+b^2=12^2\] \[b^2=144-100\] \[b^2=44\] \[b=\sqrt{44}=2\sqrt{11}\]

OpenStudy (anonymous):

I get to the point in solving the problem where you do the square root of blank = A and then i get lost.

OpenStudy (anonymous):



OpenStudy (anonymous):

I don't understand how you got that or what its for.

OpenStudy (anonymous):

now it is a matter of knowing the ratios \[\sin(\theta)=\frac{2\sqrt{11}}{12}=\frac{\sqrt{11}}{6}\] \[\cos(\theta)=\frac{10}{12}=\frac{5}{6}\] \[\tan(\theta )=\frac{2\sqrt{11}}{10}=\frac{\sqrt{11}}{5}\]

OpenStudy (anonymous):

Oh! so you simplified ?

OpenStudy (anonymous):

I'm sorry I'm terrible when it comes to math..

OpenStudy (anonymous):

 \[a^2+b^2=c^2\] if you know any two of the numbers you can find the third

OpenStudy (anonymous):

you get \[10^2+b^2=12^2\] \[100+b^2=144\] \[b^2=44\] \[b=\sqrt{44}\]

OpenStudy (anonymous):

I understand the a2+b2=c2 but I get through like maybe 3 steps and et confused.

OpenStudy (anonymous):

are those steps ok?

OpenStudy (anonymous):

Yes. It's right after that last step I don't know what to do.

OpenStudy (anonymous):

I have a dvd that explains stuff for these problems and he's saying you multiply something to get some letter and im just not sure whats going on.

OpenStudy (anonymous):

\[44=4\times 11\] \[\sqrt{44}=\sqrt{4\times 11}=\sqrt{4}\times \sqrt{11}=2\sqrt{11}\]

OpenStudy (anonymous):

it is just another way to write \(\sqrt{44}\) so you can figure out how to cancel when you take the ratios

OpenStudy (anonymous):

Ok, so you take the number (44), then do the square root of 4 and 11 since they make 44 when multiplied then how did you get 2 sqrt 11?

OpenStudy (anonymous):

i know that the square root of 4 is 2 so i replace \(\sqrt{4}\) by \(2\) i don't know what the square root of 11 is, so i just leave it as \(\sqrt{11}\)

OpenStudy (anonymous):

Oh! Okay. This is making more sense now. Thank you so much for your time/help!